# Progressive multi-fidelity learning for physical system predictions

### v1

@@ 4.1 Multi-fidelity setting
Low-fidelity input and high-fidelity output data are generated through a finite element approximation of (7) with quadratic finite elements and backward differentiation formula provided by the MATLAB library redbKIT [45]. We consider a training set computed over a short time window $T_{\text{train}}=15\text{s}<T=20\text{s}$ at $N_{\mu}=15$ Reynolds numbers equispaced over $\mathcal{P}$. High-fidelity data are computed over a mesh with 16478 triangular elements and $d_{\text{out}}=8239$ vertices, while the temporal discretization is with step size $\Delta t=5\text{ms}$. The high-fidelity output data consists of the time evolution of the velocity magnitude at the $d_{\text{out}}$ spatial degrees of freedom (dof). The input dimensions of low-fidelity data are $d_{\text{in}}^{(0)}=2,d_{\text{in}}^{(1)}=2$ (time $t$ and parameter $\mu$ at level 0, and drag and lift at level 1), and $d_{\text{in}}^{(2)}=20,d_{\text{in}}^{(3)}=577$ (spatial degrees of freedom on the outlet boundary and on a portion of the domain surrounding the obstacle; see Fig. 5). Further implementation details and training specifications are provided in A.
@@ 4.2 Results
For unseen Reynolds numbers $Re\in\{63,92\}$ and different time instances $t\in\{9\text{s},19\text{s}\}$ (with $0\text{s},15\text{s}]$ the training window), Fig. 6 shows the predicted high-fidelity velocity magnitude at each level of the progressive MF model, along with the corresponding absolute errors. At level 0, the surrogate relies solely on $(t,Re)$ as inputs and can only reproduce the average stationary behavior of the flow, missing the unsteady vortex-shedding dynamics. Incorporating drag and lift coefficients at level 1 significantly improves predictions, despite the absence of any direct measurements of the solution field. Nonetheless, non-negligible errors remain in the wake region, where nonlinear effects dominate. At level 2, additional measurements from sensors on the outlet boundary enable accurate and spatially homogeneous predictions, even beyond the training time window. Notably, the accuracy at $t=19\text{s}$ is comparable to that at $t=9\text{s}$, demonstrating the model’s capability to reliably forecast forward in time.. Remarkably, improvements are observed throughout the domain – including the cylinder wake, where no sensors are placed – highlighting the model’s ability to fuse and effectively exploit localized information. Finally, level 3, which assimilates partial-domain snapshots around the obstacle, yields only marginal gains over level 2, indicating that the outlet boundary data, combined with drag and lift measurements, already provide sufficiently rich information for the neural architecture. Table 2 reports the overall error over the full parametric, spatio-temporal test set, further supporting these conclusions.
Level 0 (input param.) Level 1 (drag/lift) Level 2 (outflow) Level 3 (partial domain)
@@
Table 2: Relative error with respect to the HF reference solution, $\mathbf{y}$, is evaluated for the MF estimates, $\hat{\mathbf{y}}^{(l)}$, at each level $l$, as $\text{err}_{\%}^{(l)}=\frac{100\%}{N_{\text{test}}}\sum_{n=1}^{N_{\text{test}}}\frac{\norm{\mathbf{y}(t_{n},\mu_{n})-\hat{\mathbf{y}}^{(l)}(t_{n},\mu_{n})}_{2}}{\norm{{\mathbf{y}}(t_{n},\mu_{n})}_{2}}$, where $N_{\text{test}}$ is the number of time-parameter combinations in the test set.
@@ 5.1 Multi-fidelity setting
The data used in this work were provided by De Vito et al. [47], collected using a gas multisensor device deployed in an Italian city. The dataset is publicly available at UCI Machine Learning Repository.
@@ 6 Conclusions
We demonstrated the flexibility and effectiveness of our method across three distinct applications: two involving simulated data (a nonlinear reaction-diffusion system and a Navier-Stokes fluid flow problem) and one real-world task (air pollution monitoring). The results consistently demonstrated that the progressive fusion of information allows the model to efficiently leverage knowledge from low-fidelity data, leading to a systematic increase in prediction accuracy and a corresponding reduction in predictive uncertainty at each level. Crucially, the framework maintains operational continuity, providing reasonable estimates even when only lower-fidelity inputs are available due to computational, budget, or data assimilation constraints. The integrated uncertainty quantification further enhances the model’s reliability, even in extrapolative regimes.
Future work will focus on enhancing the framework’s flexibility and computational efficiency. A primary goal is to develop a more robust architecture that relaxes the current requirement for a strict data hierarchy, enabling the model to handle arbitrary combinations of available or missing data modalities during inference. To improve efficiency, the computationally expensive ensemble-based uncertainty quantification could be replaced with inherent probabilistic architectures, such as Bayesian neural networks or conditional neural processes [39, 34]. Finally, the generality and effectiveness of the paradigm suggest its applicability to a broader range of scientific computing problems, including inverse problems where the multi-fidelity structure could be naturally embedded within multi-level inference schemes.
@@ Acknowledgment
PC has been supported under the JRC STEAM STM-Politecnico di Milano agreement and by the PRIN 2022 Project “Numerical approximation of uncertainty quantification problems for PDEs by multi-fidelity methods (UQ-FLY)” (No. 202222PACR), funded by the European Union - NextGenerationEU.
AM acknowledges the project “Dipartimento di Eccellenza” 2023-2027 funded by MUR, the project FAIR (Future Artificial Intelligence Research), funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, Investment 1.3, Line on Artificial Intelligence) and the Project “Reduced Order Modeling and Deep Learning for the real- time approximation of PDEs (DREAM)” (Starting Grant No. FIS00003154), funded by the Italian Science Fund (FIS) - Ministero dell’Università e della Ricerca.
@@
PC thanks Johannes Lips for fruitful discussions on the air pollution example.
Appendix A Appendix
@@
Table 4: Summary of encoder and decoder network architectures and training hyperparameters for the different applications.
This appendix provides a summary of the key hyperparameters and network architectures used for the different case studies presented in the main text. The specific choices for each application are detailed in Table 4.
@@
For the decoders, the input dimension is given by $d^{(l)}{h_{\text{tot}}}=\sum_{j\leq l}d_{h}^{(j)}$ at level $l$, while the output dimension is $d_{\text{out}}$ at every level. The decoder architectures and the number of layers/nodes are indicated under decoder type and decoder NN layers/nodes, respectively, and are the same at every level since a single high-fidelity quantity is modeled.
Moreover, Table 4 lists general NN hyperparameters used for training. The Adam optimizer was used in all cases, with learning rates and $L_{2}$ regularization coefficients tuned individually. The $\tanh$ activation function provided stable training and good performance across all applications. The regularization weights $\lambda_{\Phi}$ and $\lambda_{\Psi}$ for the encoder and decoder networks, respectively, were kept equal to 1.0 in all experiments, with the overall regularization strength $\lambda_{\text{reg}}$ controlled by the reported value.
Finally, $m$ indicates the number of independent retrainings with randomly initialized weights used to compute statistical moments for ensemble-based uncertainty quantification.

### v2

@@ 4.1 Multi-fidelity setting
Low-fidelity input and high-fidelity output data are generated through a finite element approximation of (7) with quadratic finite elements and backward differentiation formula provided by the MATLAB library redbKIT [37]. We consider a training set computed over a short time window $T_{\text{train}}=15\text{s}<T=20\text{s}$ at $N_{\mu}=15$ Reynolds numbers equispaced over $\mathcal{P}$. High-fidelity data are computed over a mesh with 16478 triangular elements and $d_{\text{out}}=8239$ vertices, while the temporal discretization is with step size $\Delta t=5\text{ms}$. The high-fidelity output data consists of the time evolution of the velocity magnitude at the $d_{\text{out}}$ spatial degrees of freedom (dof). The dimensionality of the high-fidelity output data is reduced using POD, retaining 32 modes, which capture approximately 97% of the total variance. The input dimensions of low-fidelity data are $d_{\text{in}}^{(0)}=2,d_{\text{in}}^{(1)}=2$ (time $t$ and parameter $\mu$ at level 0, and drag and lift at level 1), and $d_{\text{in}}^{(2)}=20,d_{\text{in}}^{(3)}=577$ (spatial degrees of freedom on the outlet boundary and on a portion of the domain surrounding the obstacle; see Fig. 5). Further implementation details and training specifications are provided in A.
@@ 4.2 Results
For unseen Reynolds numbers $Re\in\{63,92\}$ and different time instances $t\in\{9\text{s},19\text{s}\}$ (with $0\text{s},15\text{s}]$ the training window), Fig. 6 shows the predicted high-fidelity velocity magnitude at each level of the progressive MF model, along with the corresponding absolute errors. At level 0, the surrogate relies solely on $(t,Re)$ as inputs and can only reproduce the average stationary behavior of the flow, missing the unsteady vortex-shedding dynamics. Incorporating drag and lift coefficients at level 1 significantly improves predictions, despite the absence of any direct measurements of the solution field. Nonetheless, non-negligible errors remain in the wake region, where nonlinear effects dominate. At level 2, additional measurements from sensors on the outlet boundary enable accurate and spatially homogeneous predictions, even beyond the training time window. Notably, the accuracy at $t=19\text{s}$ is comparable to that at $t=9\text{s}$, demonstrating the model’s capability to reliably forecast forward in time.. Remarkably, improvements are observed throughout the domain – including the cylinder wake, where no sensors are placed – highlighting the model’s ability to fuse and effectively exploit localized information. Finally, level 3, which assimilates partial-domain snapshots around the obstacle, yields only marginal gains over level 2, indicating that the outlet boundary data, combined with drag and lift measurements, already provide sufficiently rich information for the neural architecture. Table 2 reports the overall error over the full parametric, spatio-temporal test set, further supporting these conclusions. Additionally, as in the reaction–diffusion example, we report the relative error on the POD expansion coefficients. The small discrepancy between the reduced-space error and the reconstructed full-field error indicates that POD truncation is negligible in this case. Consequently, the overall error is primarily driven by the progressive neural network’s ability to predict the POD coefficients, confirming that the method’s performance is governed by the surrogate model’s accuracy.
Table 2: Relative error with respect to the HF reference solution, $\mathbf{y}$, is evaluated for the MF estimates, $\hat{\mathbf{y}}^{(l)}$, at each level $l$, as $\text{err}_{\%}^{(l)}=\frac{100\%}{N_{\text{test}}}\sum_{n=1}^{N_{\text{test}}}\frac{\norm{\mathbf{y}(t_{n},\mu_{n})-\hat{\mathbf{y}}^{(l)}(t_{n},\mu_{n})}_{2}}{\norm{{\mathbf{y}}(t_{n},\mu_{n})}_{2}}$, where $N_{\text{test}}$ is the number of time-parameter combinations in the test set. Additionally, the table reports the relative error on the predicted POD expansion coefficients.
Level 0 (param.) Level 1 (drag/lift) Level 2 (outflow) Level 3 (part. domain)
@@
Rel. err. (POD coeff.) 12.3% 8.69% 6.76% 6.44%
@@ 5.1 Multi-fidelity setting
The data used in this example are extracted from the work by De Vito et al. [10] and are publicly available through the UCI Machine Learning Repository. LF inputs include chemoresistive gas sensors targeting CO and $\text{O}_{3}$, measured by the Pirelli Labs multi-sensor device, designed for continuous urban monitoring with a sensor array, data processing unit, and communication module. Two additional slots host commercial temperature and humidity sensors (levels 0 and 1, respectively), which are less prone to drift or replacement, providing robust and continuously available LF signals. CO and $\text{O}_{3}$ sensors are more sensitive to environmental conditions and aging, and may require replacement or recalibration, constituting higher-level LF inputs (levels 2 and 3) whose availability can vary over time. HF outputs correspond to benzene concentrations measured by reference-grade analyzers [10]. All measurements are preprocessed by removing invalid entries (sentinel value 200-200) and filling missing values via linear interpolation independently for each variable. Finally, all signals are rescaled to the unit interval. No additional filtering or denoising is applied, ensuring that model evaluation reflects realistic measurement noise and sensor irregularities.
@@ 6 Conclusions
We demonstrated the flexibility and effectiveness of our method across three distinct applications: two involving simulated data (a nonlinear reaction-diffusion system and a Navier-Stokes fluid flow problem) and one real-world task (air pollution monitoring). The results consistently demonstrated that the progressive fusion of information allows the model to efficiently leverage knowledge from low-fidelity data, leading to a systematic increase in prediction accuracy and a corresponding reduction in predictive uncertainty at each level. Crucially, the framework maintains operational continuity, providing reasonable estimates even when only lower-fidelity inputs are available due to computational, budget, or data assimilation constraints. The integrated uncertainty quantification further enhances the model’s reliability, even in extrapolative regimes. Regarding generalization, the air pollution application showcases strong capabilities on pure temporal forecasting, while the PDE benchmarks assess spatio-temporal generalization through temporal extrapolation and parameter interpolation. For nonlinear systems, extrapolation across parameter regimes with qualitatively different dynamics may be ill-posed for data-driven surrogates. By leveraging the progressive structure and physically meaningful low-fidelity data, the framework improves generalization in time and across parameter regimes where correlations between fidelity levels remain consistent. Conversely, for parameter regimes with fundamentally different dynamics, generalization gains may be limited, highlighting an important avenue for future study.
Moreover, future work will focus on enhancing the framework’s flexibility and computational efficiency. A primary direction is the development of multi-fidelity architectures that do not rely on a strict hierarchy of inputs. In the current progressive formulation, if an intermediate fidelity level is missing, higher-level inputs cannot be exploited, which can be limiting in multi-modal applications where a clear fidelity ordering cannot be established. The hierarchical assumption also increases computational cost: fidelity levels are added sequentially, so each additional level requires training an additional surrogate level. Early-fusion strategies combined with attention mechanisms, for instance, allow the model to dynamically weight the contribution of each available fidelity level, enabling inference with arbitrary combinations of missing or available inputs [22, 21, 16]. However, relaxing the hierarchical structure generally removes the structural guarantees against catastrophic forgetting provided by the progressive training strategy adopted here, where previously trained components are frozen and higher-fidelity information is incorporated exclusively through additive corrections. Designing flexible fusion mechanisms that retain such guarantees remains an open research challenge. Another promising direction is to improve efficiency, such that the computationally expensive ensemble-based uncertainty quantification could be replaced with inherent probabilistic architectures, such as Bayesian neural networks or conditional neural processes [14, 56].
Finally, the generality and effectiveness of the paradigm suggest its applicability to a broader range of scientific computing problems, including inverse problems where the multi-fidelity structure could be naturally embedded within multi-level inference schemes.
@@ Acknowledgment
PC has been supported under the JRC STEAM STM-Politecnico di Milano agreement and by the PRIN 2022 Project “Numerical approximation of uncertainty quantification problems for PDEs by multi-fidelity methods (UQ-FLY)” (No. 202222PACR), funded by the European Union - NextGenerationEU. PC also thanks Dr. Johannes Lips (University of Stuttgart) for fruitful discussions on the air pollution example.
AM acknowledges the project “Dipartimento di Eccellenza” 2023-2027 funded by MUR, the project FAIR (Future Artificial Intelligence Research), funded by the NextGenerationEU program within the PNRR-PE-AI scheme (M4C2, Investment 1.3, Line on Artificial Intelligence) and the Project “Reduced Order Modeling and Deep Learning for the real- time approximation of PDEs (DREAM)” (Starting Grant No. FIS00003154), funded by the Italian Science Fund (FIS) - Ministero dell’Università e della Ricerca. AM is member of the Gruppo Nazionale Calcolo Scientifico-Istituto Nazionale di Alta Matematica (GNCS-INdAM).
@@
Appendix A
@@
Offline training time per epoch
   Level 0 0.24 s 0.13 s 0.13 s
   Level 1 0.35 s 0.16 s 0.21 s
   Level 2 0.57 s 0.22 s 0.26 s
   Level 3 0.32 s 0.37 s
Online data cost (rel. to HF)
   Level 0 0% (no costs) 0% (no costs)
   Level 1 0.04% 0.02%
   Level 2 10.6% 0.27%
   Level 3 7.27%
Table 1: Summary of network architectures, training hyperparameters and computational deaitls for the different applications.
This appendix provides a summary of the implementation, key hyperparameters and network architectures used for the different case studies presented in the main text. The specific choices for each application are detailed in Table 1.
@@
For the decoders, the input dimension is given by $d^{(l)}_{h_{\text{tot}}}=\sum_{j\leq l}d_{h}^{(j)}$ at level $l$, while the output dimension is $d_{\text{out}}$ at every level. The decoder architectures and the number of layers/nodes are indicated under decoder type and decoder NN layers/nodes, respectively, and are the same at every level since a single high-fidelity quantity is modeled.
Moreover, Table 1 lists general NN hyperparameters used for training. The Adam optimizer was used in all cases, with learning rates and $L_{2}$ regularization coefficients tuned individually. All neural networks employ the $\tanh$ activation function, which is well suited for regression tasks involving time-series data, providing smooth and bounded activations that support stable training. The regularization weights $\lambda_{\Phi}$ and $\lambda_{\Psi}$ for the encoder and decoder networks, respectively, were kept equal to 1.0 in all experiments, with the overall regularization strength $\lambda_{\text{reg}}$ controlled by the reported value.
We also report, the number of independent retrainings, $m$, with randomly initialized weights used to compute statistical moments for ensemble-based uncertainty quantification. Glorot uniform initialization scheme is used for all network weights, ensuring consistent and unbiased starting points for each ensemble member. The multiple trainings are independent so they can be executed in parallel; otherwise, the computational cost scales linearly with $m$. The ensemble approach provides an implementation-agnostic estimate of epistemic uncertainty, as it does not require modifications to the training procedure or network architecture. For the Navier–Stokes example, only $m=3$ independent retrainings were performed. This number is insufficient to obtain reliable estimates of higher-order statistical moments, in particular the variance. For this reason, uncertainty measures were not reported for this case, and only ensemble-averaged predictions were shown. In large-scale or time-sensitive settings, when the computational overhead of ensemble retraining outweighs its simplicity and generality, one may instead adopt probabilistic surrogates such as Bayesian neural networks or conditional neural processes [14, 56], which provide uncertainty estimates without repeated retrainings.
Table 1 also reports indicative computational costs associated with the proposed framework. The offline training time per epoch corresponds to the average wall-clock time required to train the neural networks at each fidelity level, measured on the same hardware platform (MacBook Apple M1 8-core processor). The online data cost (relative to HF) provides an approximate estimate of the cost required to generate or acquire the input data at each fidelity level during inference, expressed as a percentage of the cost of computing a full high-fidelity solution, which is normalized to 100%. These estimates are based on the relative number of degrees of freedom of the corresponding data sources. For the reaction-diffusion and Navier-Stokes examples, Level 0 incurs zero online cost, as it only requires control parameters. For the air-pollution case study, online data costs are not reported, as the inputs consist of real sensor data and acquisition and assimilation costs are not provided.
Different normalization strategies are adopted depending on the application. For the reaction–diffusion and air-pollution examples, all inputs and outputs are rescaled to the unit interval using max-scaling applied independently to each variable. For the Navier–Stokes example, the HF output POD coefficients are instead normalized by their corresponding singular values. This ensures comparable magnitudes across modes, avoids bias toward high-energy modes during training, and improves the prediction of lower-variance yet dynamically informative components. Remaining implementation details are available in the accompanying codebase at https://github.com/ContiPaolo/Progressive-Multifidelity-NNs, ensuring full reproducibility of the results.